Choose the Correct answers: 1. A police officer spots a speeding car traveling at 20 m/s (approximately 44 mph) as it passes The officer immediately starts to drive, accelerating from rest at a rate of 5 m/s2. How far will the officer have to drive betore catchıng up to the speeding car? (a) 100 m (b) 80 m (c) 500 m (d) 16 m (e) 160 m 2. A ball is thrown with velocity of 10m/s upwards. If the ball is caught 1 m above its initial position, what is the speed of the ball when it is caught? (a) 5.3 m/s (b) 6.8 m/s (c) 9.0 m/s (d) 10 m/s (e) 22 m/s
Answer #1

1.

Using the formula x=v_xt+rac{1}{2}a_xt^2

For speeding car Vz=20 m/s and a_x=0,m/s^2

Displacement of speeding car in time tSi=20 * t

For police car v_x=0,m/s and a_x=5,m/s^2

Displacement of police car in the same time  S_2=0*t+rac{1}{2}*5*t^2

When they catch up with each other,  S1 = 20 * t = S2 = 0 * t +-*5*t2

20 * t =-*5*t

t=8,s

Displacement of police car S_2=rac{1}{2}*5*8^2=160,m

Answer is e) 160 m

2.

Using the formula v_f^2-v_i^2=2aDelta x

Initial velocity v_i=10,m/s

Acceleration Cl 9.81 m/s

Displacement Delta x=1,m

1,2-10-= 2(-9.81)(1)

Final velocity v_f=sqrt{10^2-2(9.81)}=8.965approx9.0,m/s

Answer is (c) 9.0 m/s